A mixture of air and water vapour at 1.013 bar and 16 degree C has a dew point of 5degree C.calculate the relative and specific humidity's.
1
Expert's answer
2019-07-02T08:55:27-0400

Using table for saturated vapor


We have:

Partial pressure for saturated vapor for 16 degree C

"p''_s=0.01818 (bar)"

Specific volume

"\frac{1}{\rho''}=147.2 (m^3/kg)"

From this equation

"p''=\frac{\rho''RT}{M}"

molar mass of water

"M=0.018 (kg\/mol)"

For relative humidity

"\phi=\frac{p''}{p''_s}\cdot 100%""\phi=\frac{RT}{\frac{1}{\rho''}Mp''_s}\cdot 100""\phi=\frac{8.31\cdot 289}{147.2 \cdot 0.018 \cdot 1818}\cdot 100=50 \%""E=mc^2"

Partial pressure of air we can receive from Dalton's law


"p=p'+p''""p'=p-p''"

"p'=101300-906=1.00394 (bar)"

Further, for air density from gas equation

"\rho'=\frac{p'M}{RT}"

"\rho'=\frac{101394\cdot 0.029}{8.31\cdot 289}=1.212 (kg/m^3)"

Density of moist air


"\rho=\rho'+\rho''"

Specific humidity

"SH=\frac{\rho''}{\rho'+\rho''}"

"SH=\frac{\frac{1}{147.2}}{1.212+\frac{1}{147.2}}=0.0056"